OpenStudy (anonymous):

multiplyl and simplify 5^28 * 5^10

jimthompson5910 (jim_thompson5910):

5^28 * 5^10 = 5^(28+10) = 5^(38) So \[\Large 5^{28} * 5^{10} = 5^{38}\]

OpenStudy (anonymous):

would it be 5 times 5 then add the exponents

jimthompson5910 (jim_thompson5910):

nope, just add the exponents and that sum is your final exponent

OpenStudy (anonymous):

says type in exponetial notation

jimthompson5910 (jim_thompson5910):

yep, the answer there is in exponential notation